Accountancy

One million (1,000,000) shares of stock were issued to the founders of a start-up company on January 1, 2019.   On July 1, 2019, the firm raised $500,000 from outside investors and issued 500,000 of convertible preferred stock with a conversion price of $1 per share (ROUND 1). At the time capital was raided, the investors insisted on a FIVE YEAR vesting schedule for the founders’ initial equity interest, with vesting on a monthly basis (1/60 of the founders’ interest vests each month); and weighted average anti-dilution protection in the case of a subsequent DOWN ROUND. Operations during the first eighteen months have been sluggish, and the company needs additional capital to continue operations and grow the firm’s sales and value. A second set of outside investors is willing to provide $1,300,000 in return for 2,000,000 shares of convertible preferred stock (ROUND 2). For purposes of this problem, today is July 1, 2020.

A) Given the down round, what is the new conversion price for Round 1 shares?

B) How many shares will Round 1 investors get if they convert their interest to common stock?

C) As of July 1, 2020, what percentage interests would the three parties have in the equity of the firm? (On an as if converted basis)

If no additional funding is raised, what will the equity percentages be on January 1, 2024? (Also assume as if converted basis)


Thank you..
Answer #1

Solution

Convertible Preferred Stock is that financial instruments which can be converted into equity share in future at conversion price on the choice of the shareholders.

Anti Dilution Protection is a kind of protection available to the investors to protect themselves against the dilution of their ownership in the equity of the company.

As the investors insisted on FIVE YEAR vesting schedule for the founders’ initial equity interest, with vesting on a monthly basis (1/60 of the founders’ interest vests each month)

Founder Shares in Equity on July 1, 2020, will be

= 5,00,000 *1/60*12

= 100,000 + 1,000,000

= 1,100,000 Shares

Therefore with the option of Vesting Schedule of 5 Years, Founders will get 1,00,000 shares after one year that is on July 1, 2020. and the number of shares of investors of Round 1 will be reduced by 1,00,000 after one year. Therefore Remaining Shares of Investors of Round 1 will be on July 1, 2020 :

= 5,00,000-1,00,000

=4,00,000 Shares

Existing Shareholders Structure:

Ordinary Shares: 1,000,000

Preferred Series Round 1 Shares (PMS) : 500,000

Fully Diluted Capital of Round 1 Shareholders : 500,000* $1 = $5,00,000

Price per Share of Round 1 ( PPSa) : $1

Down Round -2

Price per share of Round 2 (PPSb) : $1,300,000/2,000,000 = 0.65

Preferred Series Round 2 Shares (NS) : 2,000,000

Total Investment = $1,300,000

As the Investors of Round 1 opt for Weighted Average Anti-Dilution protection, the weighted average price of the share for Round 1 Investor will be:

Net Weighted Average Price (NBWA) for Round 1 Investors

Finance homework question answer, step 3, image 1

Therefore, with the NBWA mechanism, each Round 1 Share has been economically diluted by $0.2 (= $1- $0.8). With the NBWA formula, if the investor had subscribed shares to $0.8, the investor would have 5,00,000 shares (= $4,00,00 / $0.8), compared to the 4,00,000 acquired after Vesting Schedule of five years (5,00,000-1,00,000) = 4,00,000, so the investor will receive 1,00,000 anti-dilution shares.

    No of Shares of Round 1 investor after Anti-Dilution Protection = 4,00,000 /.8 = 5,00,000

Calculation of Percentage Interest of three parties in Equity of the firm is as follows:

Finance homework question answer, step 4, image 1

Answer: % of Founders in Equity of the Firm on July 1, 2020, is 30.55%

% of Investors of Round 1 in Equity of the Firm on July 1, 2020, is 13.89%

% of Investors of Round 2 in Equity of the Firm on July 1, 2020, is 55.56%.